OpenStudy (anonymous):

A weekly salary was increased by 8% to $237.60. What was the salary before the increase?

OpenStudy (anonymous):

.08(237.60)=19.0 237.60-19= $218.60

OpenStudy (anonymous):

But wouldnt that only be giving me 8% of the new salary and not 8% of the old salary?

OpenStudy (anonymous):

0.08*x+x=$237.60

OpenStudy (anonymous):

so: x=237,60/1,08=220

OpenStudy (anonymous):

So the old salary was $220

OpenStudy (anonymous):

ye, good job

OpenStudy (anonymous):

Ok, yay! Thank you.

OpenStudy (anonymous):

yw